solve the system of linear equations using the substitution method
2x-y=8
7x-3y=27

Answer by filboi(35)

2x-y=8; y=2x-8
7x-3y=27
7x-3(2x-8)=27
7x-6x+24=27
x=27-24
x=3
y=2(3)-8
y=-2